QUESTION 9 A compeny has the choice of spending $50,000 on either Project A or Project B. Project A will generate an additional $100,000 of revenue for the company in the first year after the ss additional revenue for the company each year for the ten years after the $50,000 expenditure is made. Based on the Payback 0,000 expenditure is made. Project 8 will generate fst Method the company 1. Project A 2. Project B QUESTION 10 If the net present value of a project is zero besed on a discount rate of 16s, then the internel rate of return is 1 , equal to 16%, o 2 less than 16%. e 3.greater than 1696. o 4 none of the above
Answer #1

Part 1:

Payback Period

Ranking

Project A

0.5 Years

I

Project B

5 years

II

Project A would be able to fully recover the initial investment in lesser period (0.5 years) compared to Project B, which needs 5 years to fully recover the initial investment. Hence, Project A is ranked the first.

  1. Computation of payback period for each project and raking:

Payback period = initial investment/annual cash inflows

Project A –

Initial investment = $50,000

Annual cash inflows = $100,000

Payback period = 50,000/100,000 = 0.5 years or 6 months

Project B –

Initial investment = $50,000

Annual cash inflows = $10,000

Payback period = 50,000/10,000 = 5 years

Part 2:

The correct option is - 1. Equal to 16%

Explanation: Internal rate of return is that discount rate that makes the net present value of a project zero. At this discount rate, the present value of estimated cash inflows from a project equal to the present value of cash outflows.

If the NPV of the project is 1 (more than zero), the cash inflows are more than the interest rate at which the future cash amounts are discounted.

If the NPV is less, that indicates that the earnings from the project are less than the interest rate at which the future cash amounts are discounted.

Since, the NPV is zero for the discount rate of 16%, the Internal Rate of Return for the given project is equal to16%.